OpenStudy (njabulo):

solve for y(y-3)/(y-2) = 2

OpenStudy (anonymous):

(y-3) = (y-2)(2) distribute the two on the right, then move the "y" variable onto one side.

OpenStudy (anonymous):

(y-3)/(y-2) = 2 ===? y-3==2y-4 ===>y=1

jhonyy9 (jhonyy9):

first of all we need suppos that y not can be never 2 because than the fraction wil be undefined